OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)? 1 2 -2 -1

OpenStudy (alexandervonhumboldt2):

(4+2)/(-4+1)=-2

OpenStudy (imtiaz7):

m(slope)=y2-y1/x2-x1 here x1=-4 and y1=4 and x2=-1 and y2=-2 then by solving we get -2

OpenStudy (anonymous):

\[\frac{-2-4}{-1--4}=-2 \]